MATRIX ISOLATION SPECTROSCOPY OF UO_{2}Cl_{2}.

\underline{CHRISTOPHER J. LUE}, JIN JIN, MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322; ,.

A search for fluorescent transitions of UCl_{4} was undertaken using samples that were isolated in solid Ar at 15K. Pulsed laser excitation was examined using the harmonics from an Nd/YAG laser (266 and 355 nm) and a dye laser operating in the 447-481 nm range. Several absorption and emission band systems were observed, with the emission spectra spanning the entire visible range. Fluorescence decay lifetimes were found to be either fast (a few hundred ns) or slow (tens of microseconds). Analysis of the vibrational progressions indicated that the carrier of these band systems was UO_{2}Cl_{2}, rather than UCl_{4}. Although present as a minor contaminant the UO_{2}Cl_{2} clearly has a much higher fluorescence quantum yield than UCl_{4}. IR spectroscopy was used to confirm the presence of UO_{2}Cl_{2} in the matrix (absorptions at 898 and 969 cm^{−1}). Assignment of the electronic transitions of UO_{2}Cl_{2} is discussed in terms of the electronic structure of the UO_{2}^{2+} sub-unit.